# The Gauss-Markov Adjunction Provides Categorical Semantics of Residuals in Supervised Learning

### v2

@@ Abstract
Denotational semantics, Adjunction, Lawvere-enriched category, AI Explicability, AI modeling.
@@ 2.3 Lawvere-enriched Category
Two Lawvere-enriched categories corresponding to systems of parameters and data will be defined later in order to describe supervised learning. The notion of metric enrichment was introduced by Lawvere [20], which demonstrated the connection between generalized metric spaces and categories. A Lawvere-enriched category $\bf C$} consists of the objects $c_{1},c_{2},\cdots\in{\rm Ob}({\bf C})$ together with Hom-objects ${\bf C}(c_{1},c_{2})={\rm Hom}_{\bf C}(c_{1},c_{2})$ [18]. These Hom-objects are to be regarded as generalized substitutes for arrows of ordinary categories, and they are objects of the base category ${\bf V}=([0,\infty],\geq,+,0)$. Each Hom-object in $\bf C$} can be identified with a distance between objects, ${\bf C}(c_{1},c_{2})=d(c_{1},c_{2})\in[0,\infty]$. In particular, the identity morphism is assigned by $d(c,c)=0$. The composition of Hom-objects is defined via the $\bf V$}-morphism ${\bf C}(c_{2},c_{3})\otimes{\bf C}(c_{1},c_{2})\to{\bf C}(c_{1},c_{3})$, which corresponds to the triangle inequality $d(c_{2},c_{3})+d(c_{1},c_{2})\geq d(c_{1},c_{3})$. A $\bf V$}-functor $F:{\bf C}\to{\bf D}$ assigns to each object $c\in{\rm Ob}({\bf C})$ an object $Fc\in{\rm Ob}({\bf D})$, and is determined by the collection of all morphisms $F_{c_{1},c_{2}}:{\bf C}(c_{1},c_{2})\to{\bf D}(Fc_{1},Fc_{2})$, which coherently preserve identities and compositions across the Hom-objects.
@@ 3 Categories and Functors for Regression
Fix an arbitrary matrix $X\in\mathbb{R}^{n\times m}$, and let $G\in\mathbb{R}^{m\times n}$ denote the Moore-Penrose inverse of $X$. Fix an arbitrary vector $b\in\mathbb{R}^{n}$, on which we may impose the constraint $b=b_{\perp}\in{\rm ker}(X^{\top})$ when required. Under this constraint, it follows that $X^{\top}b_{\perp}=0$ and $Gb_{\perp}=0$, since ${\rm ker}(G)={\rm ker}(X^{\top})$.
@@ 3.2 Affine Forward Functor and Gauss-Markov Functor
In this section, we define two functors: $\mathcal{F}:{\bf Prm}\to{\bf Data}$ and $\mathcal{G}:{\bf Data}\to{\bf Prm}$. These functors are induced by the fixed matrix and vector, $X$ and $b$.
@@
While $\mathcal{F}$ may be defined for arbitrary $b$, in cases where it is useful to emphasize a specific choice $b$, we denote it with a subscript as $\mathcal{F}_{b}$.
@@ 4 Gauss-Markov Adjunction
For the affine forward functor $\mathcal{F}$ and the Gauss-Markov functor $\mathcal{G}$, there exists a $\bf V$}-natural transformation
$\Phi:{\rm Hom}_{\bf Data}(\mathcal{F}a,y)\to{\rm Hom}_{\bf Prm}(a,\mathcal{G}y).$ (7)
@@
There exists a $\bf V$}-natural transformation
${\rm Hom}_{\bf Prm}(a,\mathcal{G}\mathcal{F}a)=\varphi(a,\mathcal{G}\mathcal{F}a)=:\omega_{a},$ (9)
which we refer to as a pre-unit, with the following property. For any objects $a\in{\rm Ob}({\bf Prm})$ and $y\in{\rm Ob}({\bf Data})$, and for a Hom-object $\varphi(a,\mathcal{G}y)$, there exists a Hom-object $\delta(\mathcal{F}a,y)$ such that
$\varphi(\mathcal{G}\mathcal{F}a,\mathcal{G}y)+\omega_{a}\;\geq\;\varphi(a,\mathcal{G}y),$ (10)
which implies
${\rm Hom}_{\bf Prm}(\mathcal{G}\mathcal{F}a,\mathcal{G}y)\,\otimes\,{\rm Hom}_{\bf Prm}(a,\mathcal{G}\mathcal{F}a)\;\longrightarrow\;{\rm Hom}_{\bf Prm}(a,\mathcal{G}y).$ (11)
In particular, for $\mathcal{F}_{b_{\perp}}$ constrained by $b=\,^{\forall}b_{\perp}\in{\rm ker}(X^{\top})$, the equality in (10) holds and $\delta(\mathcal{F}a,y)$ is uniquely determined for any $a$, $y$, and $\varphi(a,\mathcal{G}y)$; that is,
@@
This is the universal mapping property of the unit $\omega_{a,b_{\perp}}$. $\square$
@@
First, a $\bf V$}-natural transformation $\omega_{a}=\varphi(a,\mathcal{GF}a)$ is defined as a family of Hom-objects $\omega=\{\omega_{a}\}_{a\in\rm{Ob}({\bf Prm})}$ such that for every object $a$ and every Hom-object $\varphi(a,a^{\prime})$ in $\bf Prm$}, the following diagram commutes:
[…]
This definition guarantees the naturality of $\omega$. From the definitions of the functors and the Hom-objects, we obtain $\varphi(a,\mathcal{G}y)=\|X(Gy-a)\|$, $\mathcal{G}:\delta(\mathcal{F}a,y)\to\varphi(\mathcal{GF}a,\mathcal{G}y)=\|X(Gy-GXa-Gb)\|=\|X(Gy-a)-XGb\|$, and $\omega_{a}=\varphi(a,\mathcal{GF}a)=\|X\{G(Xa+b)-a\}\|=\|Xa+XGb-Xa\|=\|XGb\|$. From the triangle inequality $\|X(Gy-a)-XGb\|+\|XGb\|\geq\|X(Gy-a)\|$, we obtain that the inequality (10) holds. Moreover, if $b=b_{\perp}$, then $XGb=0$, and hence equation (12) follows. In this case, the pre-unit $\omega_{a}$ becomes the genuine unit $\omega_{a,b_{\perp}}$ of the adjunction. $\square$
In summary, Proposition (GM-2) is captured by the following commutative diagram:
[…]
The triangle in the above diagram (15), represented by equation (12), and the natural transformation of the Gauss-Markov adjunction (8) are related by
$\Phi_{\rm GM}:{\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}a,y)$ $\xrightarrow{\qquad\mathcal{G}\qquad}$ ${\rm Hom}_{\bf Prm}(\mathcal{G}\mathcal{F}_{b_{\perp}}a,\mathcal{G}y)$ (16)
$\xrightarrow{\;(-)\,\otimes\,{\rm Hom}_{\bf Prm}(a,\mathcal{G}\mathcal{F}_{b_{\perp}}a)\;}$ ${\rm Hom}_{\bf Prm}(a,\mathcal{G}y),$
where the unit $\omega_{a,b_{\perp}}$ appears in the second $\bf V$}-morphism, $(-)\,\otimes\,{\rm Hom}_{\bf Prm}(a,\mathcal{G}\mathcal{F}_{b_{\perp}}a)$.
@@
There exists a $\bf V$}-natural transformation
${\rm Hom}_{\bf Data}(\mathcal{FG}y,y)=\varphi(\mathcal{FG}y,y)=:\varepsilon_{y},$ (17)
which we refer to as a pre-counit, with the following property. For any objects $a\in{\rm Ob}({\bf Prm})$ and $y\in{\rm Ob}({\bf Data})$, and for a Hom-object $\delta(\mathcal{F}a,y)$, there exists a Hom-object $\varphi(a,\mathcal{G}y)$ such that
$\varepsilon_{y}+\delta(\mathcal{F}a,\mathcal{FG}y)\;\geq\;\delta(\mathcal{F}a,y),$ (18)
which implies
${\rm Hom}_{\bf Data}(\mathcal{FG}y,y)\,\otimes\,{\rm Hom}_{\bf Data}(\mathcal{F}a,\mathcal{FG}y)\;\longrightarrow\;{\rm Hom}_{\bf Data}(\mathcal{F}a,y).$ (19)
In particular, for $\mathcal{F}_{b_{\perp}}$, the equality in (18) holds and $\varphi(a,\mathcal{G}y)$ is uniquely determined for any $a$, $y$, and $\delta(\mathcal{F}a,y)$; that is,
@@
${\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}\mathcal{G}y,y)\,\otimes\,{\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}a,\mathcal{F}_{b_{\perp}}\mathcal{G}y)\;\longrightarrow\;{\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}a,y).$ (21)
This is the universal mapping property of the counit $\varepsilon_{y,b_{\perp}}$. $\square$
@@
From the definitions of the functors and the Hom-objects, we obtain $\delta(\mathcal{F}a,y)=\|XG(y-Xa-b)\|=\|X(Gy-a)-XGb\|$, $\mathcal{F}:\varphi(a,\mathcal{G}a)\to\delta(\mathcal{F}a,\mathcal{FG}y)=\|XG\left[\{X(Gy)+b\}-(Xa+b)\right]\|=\|X(Gy-a)\|$, and $\varepsilon_{y}=(\mathcal{FG}y,y)=\|XG(y-XGy-b)\|=\|-XGb\|$. From the triangle inequality $\|-XGb\|+\|X(Gy-a)\|\geq\|X(Gy-a)-XGb\|$, we obtain that the inequality (18) holds. Moreover, if $b=b_{\perp}$, then $XGb=0$, and hence equation (20) follows. In this case, the pre-counit $\varepsilon_{y}$ becomes the genuine unit $\varepsilon_{y,b_{\perp}}$ of the adjunction. $\square$
The proposition (GM-3) is summarized by the following diagram:
[…]
The natural transformation of the Gauss-Markov adjunction in Propositions (GM-1) and (GM-3) are related by
$\Phi_{\rm GM}:{\rm Hom}_{\bf Prm}(a,\mathcal{G}y)$ $\xrightarrow{\qquad\mathcal{F}_{b_{\perp}}\qquad}$ ${\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}a,\mathcal{F}_{b_{\perp}}\mathcal{G}y)$ (23)
$\xrightarrow{\;{\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}\mathcal{G}y,y)\,\otimes\,(-)\;}$ ${\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}a,y),$
where the counit $\varepsilon_{y}$ appears in the second $\bf V$}-morphism, ${\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}\mathcal{G}y,y)\otimes(-)$.
Consequently, the above propositions on the adjunction $\mathcal{F}_{b_{\perp}}\dashv\mathcal{G}$ are summarized in the following diagram:
${{\bf Data}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\mathcal{F}}$${{\bf Prm}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{G}}$${\qquad\Downarrow\;b=b_{\perp}}$${{\bf Data}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{F}_{b_{\perp}}}\textstyle{\perp}$${{\bf Prm}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{G}}$ (24)

### v3

@@ Abstract
Adjunction, Lawvere-enriched category, Lax natural transformation, Denotational semantics, AI Explicability.
@@ 2.1 Regression
2.2 Calibration Parameter
The multiple regression model can formally be extended by adding an arbitrary parameter vector $b\in\mathbb{R}^{n}$, resulting in the expression $y=Xa+b+r$. Note that this $b$ is neither an intercept term, which appears as the first component of the parameter vector $a$, nor a residual term $r$. Rather, the parameter $b$ represents an explicit calibration term. Mathematically, $b$ acts as a translation in the affine mapping $\mathcal{F}(a)=Xa+b$. Since this merely shifts the origin of $y$, and the expression reduces to $z=Xa+r$ by defining z=ybz=y-b, it may seem that treating $b$ explicitly is unnecessary.
However, in the categorical formulation of regression developed in the following sections, the inclusion or omission of the calibration term $b$ has a distinct impact on the structure of the corresponding diagrams. This $b$ appears in the categorical notion of a natural transformation and, in particular, serves to make the unit and counit explicit. It helps to distinguish functorial operations on objects, morphisms, and functors, thereby revealing a hierarchical semantics that remains implicit in conventional linear-algebraic regression. If $b$ is disregarded from the outset, the unit and counit become invisible, making them difficult to identify and trace within the structure of the adjunction. Consequently, retaining $b$ clarifies computations involving functors and facilitates the tracking of adjunction proofs, thereby ensuring categorical consistency.
@@ 2.4 Lawvere-enriched Category
Two Lawvere-enriched categories corresponding to systems of parameters and data will be defined later in order to describe supervised learning. The notion of metric enrichment was introduced by Lawvere [20], which demonstrated the connection between generalized metric spaces and categories. A Lawvere-enriched category $\bf C$} consists of the objects $c_{1},c_{2},\cdots\in{\rm Ob}({\bf C})$ together with Hom-objects ${\bf C}(c_{1},c_{2})={\rm Hom}_{\bf C}(c_{1},c_{2})$. These Hom-objects are to be regarded as generalized substitutes for arrows of ordinary categories, and they are objects of the base category ${\bf V}=([0,\infty],\geq,+,0)$. Each Hom-object in $\bf C$} can be identified with a distance between objects, ${\bf C}(c_{1},c_{2})=d(c_{1},c_{2})\in[0,\infty]$. In particular, the identity morphism is assigned by $d(c,c)=0$. The composition of Hom-objects is defined via the $\bf V$}-morphism ${\bf C}(c_{2},c_{3})\otimes{\bf C}(c_{1},c_{2})\to{\bf C}(c_{1},c_{3})$, which corresponds to the triangle inequality $d(c_{2},c_{3})+d(c_{1},c_{2})\geq d(c_{1},c_{3})$. A $\bf V$}-functor $T:{\bf C}\to{\bf D}$ assigns to each object $c\in{\rm Ob}({\bf C})$ an object $Tc\in{\rm Ob}({\bf D})$, and is determined by the collection of all morphisms $T_{c_{1},c_{2}}:{\bf C}(c_{1},c_{2})\to{\bf D}(Tc_{1},Tc_{2})$, which coherently preserve identities and compositions across the Hom-objects.
Kelly [18] defines a monoidal category ${\bf V}=({\bf V}_{0},\otimes,I,a,l,r)$ equipped with natural isomorphisms $a_{c_{1}c_{2}c_{3}}:(c_{1}\otimes c_{2})\otimes c_{3}\cong c_{1}\otimes(c_{2}\otimes c_{3})$, $l_{c}:I\otimes c\cong c$, $r_{c}:c\otimes I\cong c$, together with an identity element $\iota_{c}:I\to{\bf C}(c,c)$ for each object $c$. In the case of a Lawvere-enriched category, these correspond to ${\bf V}_{0}=[0,\infty]$, $\otimes=+$, and $I=0$. Under this definition, a $\bf V$}-natural transformation can be formulated as follows. For $\bf V$}-functors $T,S:{\bf C}\to{\bf D}$, a $\bf V$}-natural transformation $\alpha:T\to S$ is a family of components
$\alpha_{c}:I\to{\bf D}(Tc,Sc)$ (1)
for each object $c$, satisfying the $\bf V$}-naturality condition, which is expressed by the commutativity of the following diagram:
${I\otimes{\bf C}(c_{1},c_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{c_{2}}\otimes T\qquad}$${{\bf D}(Tc_{2},Sc_{2})\otimes{\bf D}(Tc_{1},Tc_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{composition}$${{\bf C}(c_{1},c_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{l^{-1}}\scriptstyle{\cong}$$\scriptstyle{r^{-1}}$${{\bf D}(Tc_{1},Sc_{2})}$${{\bf C}(c_{1},c_{2})\otimes I\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{S\otimes\alpha_{c_{1}}\qquad}$${{\bf D}(Sc_{1},Sc_{2})\otimes{\bf D}(Tc_{1},Sc_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\;.}$$\scriptstyle{composition}$ (2)
Applying the definition of a $\bf V$}-natural transformation (1) to the case of Lawvere-enriched categories, we obtain $\alpha_{c}:0\to{\bf D}(Tc,Sc)$. This implies $0\geq{\bf D}(Tc,Sc)=d(Tc,Sc)\geq 0$, and therefore ${\bf D}(Tc,Sc)=0$, which represents a trivial Hom-object. Note that such degeneracy is a general feature of metric enrichment.
In Section 3.3, we introduce the $\Lambda$-natural transformation, defined as a lax $\bf V$}-natural transformation, in order to resolve the degeneracy and to reveal the hidden categorical structure of regression.
@@ 3 Categories and Functors for Regression
Fix an arbitrary matrix $X\in\mathbb{R}^{n\times m}$, and let $G\in\mathbb{R}^{m\times n}$ denote the Moore-Penrose inverse of $X$. Select an arbitrary vector $b\in\mathbb{R}^{n}$, on which we may impose the constraint $b=b_{\perp}\in{\rm ker}(X^{\top})$ when required. Under this constraint, it follows that $X^{\top}b_{\perp}=0$ and $Gb_{\perp}=0$, since ${\rm ker}(G)={\rm ker}(X^{\top})$. We refer to $b$ as the calibration parameter.
@@ 3.2 Affine Forward Functor and Gauss-Markov Functor
In this section, we define two functors: $\mathcal{F}:{\bf Prm}\to{\bf Data}$ and $\mathcal{G}:{\bf Data}\to{\bf Prm}$. These functors are determined by the fixed matrix $X$ and the calibration parameter $b$. While $\mathcal{F}$ may be defined for an arbitrary $b$, in cases where it is useful to emphasize a specific choice of $b$, we denote it with a subscript as $\mathcal{F}_{b}$. Accordingly, $\{\mathcal{F}_{b}\mid b\in\mathbb{R}^{n}\}$ denotes the family of functors parameterized by $b$.
@@
3.3 $\Lambda$-Natural Transformation
As shown in Section 2.4, a $\bf V$}-natural transformation (1) generally degenerates to 0 in the case of Lawvere-enriched categories. In this section, we introduce the $\Lambda$-natural transformation, defined as a lax $\bf V$}-natural transformation, in order to resolve this degeneracy and to reveal the hidden categorical structure of regression.
First, we define the following specific Hom-object,
$\Lambda:=\|XGb\|.$ (9)
This equation means that $\Lambda$ is a canonical lifting of the calibration parameter $b$, from a mere parameter vector to a Hom-object. From the affine forward functors, $\mathcal{F}_{0}:a\to Xa$ and $\mathcal{F}_{b}:a\to Xa+b$, for any $a\in{\rm Ob}({\bf Prm})$, we obtain the following equation,
${\rm Hom}_{\bf Data}(\mathcal{F}_{0}a,\mathcal{F}_{b}a)=\|XG\{(Xa+b)-Xa\}\|=\|XGb\|=\Lambda.$ (10)
Therefore, $\Lambda$ can be interpreted as the internal metric deviation between the linear and affine components of $\mathcal{F}$. It has a desirable property that is preserved under the mapping by the Gauss-Markov functor $\mathcal{G}$:
${\rm Hom}_{\bf Prm}(\mathcal{GF}_{0}a,\mathcal{GF}_{b}a)=\|X\{G(Xa+b)-GXa\}\|=\|XGb\|=\Lambda.$ (11)
Therefore, the following diagram commutes:
[…]
If $b=b_{\perp}$, then $\Lambda_{b_{\perp}}=\|XGb_{\perp}\|=0=I$, and the above diagram degenerates.
Under this definition, a $\Lambda$-natural transformation can be formulated as follows. For affine forward functors, $\mathcal{F}_{1}:a\to Xa+b_{1}$ and $\mathcal{F}_{2}:a\to Xa+b_{2}$, a $\Lambda$-natural transformation $\lambda:\mathcal{F}_{1}\to\mathcal{F}_{2}$ is a family of components
$\lambda_{a}:\Lambda\to\delta(\mathcal{F}_{1}a,\mathcal{F}_{2}a)=\|XG(b_{2}-b_{1})\|$ (13)
for each object $a$, satisfying the $\Lambda$-naturality condition, which is expressed by the commutativity of the following diagram:
${\Lambda\otimes\varphi(a_{1},a_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda_{a_{2}}\otimes\mathcal{F}_{1}\qquad}$${\delta(\mathcal{F}_{1}a_{2},\mathcal{F}_{2}a_{2})\otimes\delta(\mathcal{F}_{1}a_{1},\mathcal{F}_{1}a_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{composition}$${\delta(\mathcal{F}_{1}a_{1},\mathcal{F}_{2}a_{2})}$${\varphi(a_{1},a_{2})\otimes\Lambda\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\mathcal{F}_{2}\otimes\lambda_{a_{1}}\qquad}$${\delta(\mathcal{F}_{2}a_{1},\mathcal{F}_{2}a_{2})\otimes\delta(\mathcal{F}_{1}a_{1},\mathcal{F}_{2}a_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\;.}$$\scriptstyle{composition}$ (14)
Equations (13) and (14) together define the $\Lambda$-natural transformation as a lax $\bf V$}-natural transformation, inspired by (1) and (2). If $b\geq b_{2}-b_{1}$, then the $\Lambda$-natural transformation $\lambda_{a}$ exists, thus avoiding the degeneracy inherent in the $\bf V$}-natural transformation of Lawvere-enriched categories.
@@ 4 Gauss-Markov Adjunction
For the affine forward functor $\mathcal{F}$ and the Gauss-Markov functor $\mathcal{G}$, there exists a $\Lambda$-natural transformation
$\Phi_{\Lambda}:{\rm Hom}_{\bf Data}(\mathcal{F}a,y)\to{\rm Hom}_{\bf Prm}(a,\mathcal{G}y).$ (15)
@@
There exists a $\Lambda$-natural transformation $\bar{\omega}:\mathcal{I}\to\mathcal{GF}$,
$\bar{\omega}_{a}:\Lambda\to{\rm Hom}_{\bf Prm}(a,\mathcal{G}\mathcal{F}a)=\varphi(a,\mathcal{GF}a)=\omega_{a},$ (17)
which we refer to as a $\Lambda$-unit, with the following property. For any objects $a\in{\rm Ob}({\bf Prm})$ and $y\in{\rm Ob}({\bf Data})$, and for a Hom-object $\varphi(a,\mathcal{G}y)$, there exists a Hom-object $\delta(\mathcal{F}a,y)$ such that the following diagram commutes:
${\Lambda\otimes\delta(\mathcal{F}a,y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\bar{\omega}_{a}\otimes\mathcal{G}\qquad}$${\varphi(a,\mathcal{GF}a)\otimes\varphi(\mathcal{GF}a,\mathcal{G}y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{composition}$${\varphi(a,\mathcal{G}y)}$${\delta(\mathcal{F}a,y)\otimes\Lambda\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\mathcal{G}\otimes\bar{\omega}_{a}\qquad}$${\varphi(\mathcal{GF}a,\mathcal{G}y)\otimes\varphi(a,\mathcal{GF}a)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\;,}$$\scriptstyle{composition}$ (18)
where the right vertical arrow implies
$\varphi(\mathcal{G}\mathcal{F}a,\mathcal{G}y)+\omega_{a}\;\geq\;\varphi(a,\mathcal{G}y).$ (19)
In particular, if we select $b=\,^{\forall}b_{\perp}\in{\rm ker}(X^{\top})$, then $\omega_{a,b_{\perp}}=0$, $\Lambda_{b_{\perp}}=I=0$, the equality in (19) holds, and $\delta(\mathcal{F}a,y)$ is uniquely determined for any $a$, $y$, and $\varphi(a,\mathcal{G}y)$; that is,
@@
This is the universal mapping property of the unit $\omega_{a,b_{\perp}}$, in the sense of $\bf V$}-natural transformation. $\square$
@@
From the definitions of the functors and the Hom-objects, we obtain $\varphi(a,\mathcal{G}y)=\|X(Gy-a)\|$, $\mathcal{G}:\delta(\mathcal{F}a,y)\to\varphi(\mathcal{GF}a,\mathcal{G}y)=\|X(Gy-GXa-Gb)\|=\|X(Gy-a)-XGb\|$, and $\omega_{a}=\varphi(a,\mathcal{GF}a)=\|X\{G(Xa+b)-a\}\|=\|Xa+XGb-Xa\|=\|XGb\|$. From the triangle inequality $\|X(Gy-a)-XGb\|+\|XGb\|\geq\|X(Gy-a)\|$, we obtain that the inequality (19) holds. Moreover, if $b=b_{\perp}$, then $XGb=0$, and hence equation (20) follows. In this case, the $\Lambda$-unit $\omega_{a}$ becomes the genuine unit $\omega_{a,b_{\perp}}=0$ of the adjunction in the sense of $\bf V$}-natural transformation. $\square$
The Gauss-Markov adjunction expressed in Proposition (GM-2) is summarized by the following diagram:
${I\otimes\delta(\mathcal{F}_{b_{\perp}}a,y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\bar{\omega}_{a,b_{\perp}}\otimes\mathcal{G}\qquad}$${\varphi(a,\mathcal{GF}_{b_{\perp}}a)\otimes\varphi(\mathcal{GF}_{b_{\perp}}a,\mathcal{G}y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\delta(\mathcal{F}_{b_{\perp}}a,y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\varphi(a,\mathcal{G}y)}$${\delta(\mathcal{F}_{b_{\perp}}a,y)\otimes I\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\mathcal{G}\otimes\bar{\omega}_{a,b_{\perp}}\qquad}$${\varphi(\mathcal{GF}_{b_{\perp}}a,\mathcal{G}y)\otimes\varphi(a,\mathcal{GF}_{b_{\perp}}a)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\;.}\scriptstyle{\cong}$ (22)
@@
There exists a $\Lambda$-natural transformation $\bar{\varepsilon}:\mathcal{FG}\to\mathcal{I}$,
$\bar{\varepsilon}_{y}:\Lambda\to{\rm Hom}_{\bf Data}(\mathcal{FG}y,y)=\delta(\mathcal{FG}y,y)=\varepsilon_{y},$ (23)
which we refer to as a $\Lambda$-counit, with the following property. For any objects $a\in{\rm Ob}({\bf Prm})$ and $y\in{\rm Ob}({\bf Data})$, and for a Hom-object $\delta(\mathcal{F}a,y)$, there exists a Hom-object $\varphi(a,\mathcal{G}y)$ such that the following diagram commutes:
${\Lambda\otimes\varphi(a,\mathcal{G}y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\bar{\varepsilon}_{y}\otimes\mathcal{F}\qquad}$${\delta(\mathcal{FG}y,y)\otimes\delta(\mathcal{F}a,\mathcal{FG}y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{composition}$${\delta(\mathcal{F}a,y)}$${\varphi(a,\mathcal{G}y)\otimes\Lambda\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\mathcal{F}\otimes\bar{\varepsilon}_{y}\qquad}$${\delta(\mathcal{F}a,\mathcal{FG}y)\otimes\delta(\mathcal{FG}y,y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\;,}$$\scriptstyle{composition}$ (24)
where the right vertical arrow implies
$\varepsilon_{y}+\delta(\mathcal{F}a,\mathcal{FG}y)\;\geq\;\delta(\mathcal{F}a,y).$ (25)
In particular, if we select $b=\,^{\forall}b_{\perp}\in{\rm ker}(X^{\top})$, then $\varepsilon_{y,b_{\perp}}=0$, $\Lambda_{b_{\perp}}=I=0$, the equality in (25) holds, and $\varphi(a,\mathcal{G}y)$ is uniquely determined for any $a$, $y$, and $\delta(\mathcal{F}a,y)$; that is,
@@
${\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}\mathcal{G}y,y)\,\otimes\,{\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}a,\mathcal{F}_{b_{\perp}}\mathcal{G}y)\;\xrightarrow{\;\cong\;}\;{\rm Hom}_{\bf Data}(\mathcal{F}_{b_{\perp}}a,y).$ (27)
This is the universal mapping property of the counit $\varepsilon_{y,b_{\perp}}$, in the sense of $\bf V$}-natural transformation. $\square$
@@
From the definitions of the functors and the Hom-objects, we obtain $\delta(\mathcal{F}a,y)=\|XG(y-Xa-b)\|=\|X(Gy-a)-XGb\|$, $\mathcal{F}:\varphi(a,\mathcal{G}y)\to\delta(\mathcal{F}a,\mathcal{FG}y)=\|XG\left[\{X(Gy)+b\}-(Xa+b)\right]\|=\|X(Gy-a)\|$, and $\varepsilon_{y}=(\mathcal{FG}y,y)=\|XG(y-XGy-b)\|=\|-XGb\|$. From the triangle inequality $\|-XGb\|+\|X(Gy-a)\|\geq\|X(Gy-a)-XGb\|$, we obtain that the inequality (25) holds. Moreover, if $b=b_{\perp}$, then $XGb=0$, and hence equation (26) follows. In this case, the $\Lambda$-counit $\varepsilon_{y}$ becomes the genuine unit $\varepsilon_{y,b_{\perp}}=0$ of the adjunction in the sense of $\bf V$}-natural transformation. $\square$
The Gauss-Markov adjunction expressed in Proposition (GM-3) is summarized by the following diagram:
${I\otimes\varphi(a,\mathcal{G}y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\bar{\varepsilon}_{y,b_{\perp}}\otimes\mathcal{F}_{b_{\perp}}\qquad}$${\delta(\mathcal{F}_{b_{\perp}}\mathcal{G}y,y)\otimes\delta(\mathcal{F}_{b_{\perp}}a,\mathcal{F}_{b_{\perp}}\mathcal{G}y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\varphi(a,\mathcal{G}y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\delta(\mathcal{F}_{b_{\perp}}a,y)}$${\varphi(a,\mathcal{G}y)\otimes I\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\mathcal{F}_{b_{\perp}}\otimes\bar{\varepsilon}_{y,b_{\perp}}\qquad}$${\delta(\mathcal{F}_{b_{\perp}}a,\mathcal{F}_{b_{\perp}}\mathcal{G}y)\otimes\delta(\mathcal{F}_{b_{\perp}}\mathcal{G}y,y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\;.}\scriptstyle{\cong}$ (28)